# Quark-versus-gluon tagging in CMS Open Data with CWoLa and TopicFlow

### v1

@@ II.2 Jet topics
The jet-topics framework [33, 34, 31] provides a method for extracting these fractions. In addition to sample independence of $p_{Q}$ and $p_{G}$ this method requires that one can measure an observable in which the distributions are mutually irreducible. This further assumption is equivalent to requiring that the support of $p_{Q}$ is not a subset of the support of $p_{G}$ and vice-versa. Under these conditions, Eq. 2 can be rewritten in terms of reducibility factors, defined as
[…]
@@
Since the $\kappa$ factors are obtainable from samples of $M_{1}$ and $M_{2}$, their measurement provides a direct estimate for the quark fractions $f_{1},f_{2}$ and pure distributions that can be performed on data. Importantly, the extracted fractions apply to any observable, not only the $x$ with which the reducibility factors were determined. A number of methods for performing the minimisation of Eq. 3 are outlined in [31].
@@
Of course, categories defined in this manner are no longer completely data-driven. An alternative method for extracting mixture fractions in reducible observables is described in [79].
II.3 TopicFlow
In previous work, some of us developed TopicFlow [42], with which one can train a generative model for the topic distributions $p_{Q}$ and $p_{G}$ given estimates of the quark fractions of two mixtures. The loss function of a deep generative model usually takes the form of an expectation value for some function $L$ (e.g. the negative log-likelihood in the case of normalising flows) over samples from the target distribution. While one cannot directly sample pure quark or gluon jets from experimental data, we can use Eq. 5 to rewrite the loss in terms of expectations over mixtures $M_{i}$. Taking the quark distribution as an example, one has
@@ II.3 TopicFlow
In practice, training models with a loss that balances opposite-sign terms can harm the stability of the optimisation. In the case of the loss in Eq. 7, the network may discover that it can ‘cheat’ by ignoring the first term and trying to achieve poor loss on $M_{2}$, which is often trivial. In this work, we use an alternative loss that jointly trains quark and gluon topic models based on Eq. 1. Specifically, given a loss function $L$ that depends on the likelihood of the data under the model, we minimise [^3] For maximum likelihood training, the function $L(x)$ is simply $-\log{x}$. One can of course consider other objective functions including those that depend on the likelihood only implicitly.
@@
While it is straightforward to apply Eq. 5 to histograms of two mixed datasets, generative models offer a number of benefits compared to binned approaches, particularly when $x$ is a high-dimensional observable. Firstly, generative models allow oversampling to smooth statistics over small datasets. In some cases, subject to the available data and the inductive bias of the model, deep generative models have been observed to contain more information than the data on which they were trained [80, 81]. A second advantage is the possibility of conditioning the generative model on auxiliary inputs such that distributions can be interpolated between measured points (e.g. [82, 83]). The use case we will explore in this work, however, is generative classification where one constructs a quark/gluon likelihood ratio using individually modelled distributions.
In [42], TopicFlow was applied only to simulated data. However it is compatible with real data since the underlying model (a normalising flow) is unsupervised. This study therefore represents the first application of the method to real jets.
III Training procedure and network architecture
@@
Figure 3: Data/MC comparison for two Energy Flow Polynomials in LABEL:sub@efp_1_dijet, LABEL:sub@efp_2_dijet dijet events and LABEL:sub@efp_1_zjet, LABEL:sub@efp_2_zjet $Z+$jet events.
@@
Fully-supervised Quarks and gluons from QCD MC
@@
As a representation for the jets, we make use of Energy Flow Polynomials (EFPs) [84]. EFPs form an over-complete basis for infrared and collinear safe (IRC) jet observables and are thus suited for use with real data, since physical detectors have finite energy/angle resolution. It was also observed in [78] that IRC-safe observables exhibit smaller sample dependence.
where $M$ is the number of jet constituents, $z_{a}=p_{T,a}/p_{T,J}$ is the $p_{T}$ fraction of constituent $a$ relative to the jet, $\theta_{ab}$ is the separation of $a$ and $b$ in the rapidity-azimuth plane, and $\beta$ is a parameter. We consider connected EFPs with $\beta=\frac{1}{2}$ and number of edges $d\in[1,3\,]$. Figure 3 shows histograms for two of the eight total EFPs in the basis, corresponding to the graphs
@@
The distributions exhibit small Data-MC discrepancies, in particular systematically smaller EFP values in DY simulation compared to the DoubleMu primary dataset.
@@
Figure 4: Histograms of the first two Energy Flow Polynomials (top row) and classifier predictions (bottom row) for quarks and gluons in dijet and $Z+$jet events from CMS simulation. The agreement among similar colours illustrates approximate sample independence in the datasets.
@@
For TopicFlow, we train continuous normalising flows with the Flow Matching objective of [85]. To adapt this loss according to Eq. 8, we replace the single regression network with a linear combination of ‘quark’ and ‘gluon’ networks as $v_{\theta}\to f_{i}v_{\theta_{Q}}+(1-f_{i})v_{\theta_{G}}$. [^5] Interestingly, for two vector fields $v_{Q}$ and $v_{G}$ respectively generating probability paths $p_{t,Q}$ and $p_{t,G}$ it does not follow that the combination $fv_{Q}+(1-f)v_{G}$ generates $fp_{t,Q}+(1-f)p_{t,G}$. Despite this, we find empirically that models trained in this way learn the expected topic distributions. For each regression network we use an MLP consisting of three 512-node layers with GELU activations and a skip connection between the first and third layer. Dropout is applied after each layer at a rate of 0.2. We use the Adam optimiser with an initial learning rate of $10^{-3}$ which is decayed by a factor 10 if there is no improvement for 20 epochs. Training is halted after 200 epochs, or once the learning rate has decayed below $10^{-6}$.
IV Results
IV.1 Extracted quark fractions
@@
Here we investigate the quark/gluon content of each of the datasets using the jet topics. It is first necessary, however to assess whether or not the key assumptions of the framework are valid for our datasets. To validate the assumption of sample independence, we plot in Figure 4 the MC-matched quark and gluon distributions of two EFPs and the outputs of one of each type of classifier, both for DY and QCD simulation. In each observable, there are small but noticeable differences between the different MC datasets. However, the general shape of the distributions agree to a good approximation, with deviations mostly residing in the tails. In the EFP distributions, shown in the top row, the quark distributions are more similar across the $Z+$jet and dijet samples compared to the gluon distributions. Interestingly, when considering the outputs of the two CWoLa models in Figs. 4c and 4d (which depend only on the 8 EFPs) the opposite trend is evident. Specifically, both of these models assign slightly higher scores to DY MC quarks than QCD MC quarks, as indicated by the uptick of purple squares in the DY/QCD ratio. The predictions of the CWoLa models exhibit the greatest sample dependence of the observables we consider, with deviations up to 25% in the worst regions. Meanwhile, the fully-supervised network has outputs that are relatively consistent across the MC samples.
@@
A fully fledged measurement of any quantity requires a robust estimate of systematic uncertainties. In this essentially exploratory paper we have not attempted an analysis of these. All uncertainties presented are statistical. Consequently we refer to any quantities we derive in this paper as estimates and not as measurements.
@@
To extract reducibility factors and quark fractions in each dataset, we train classifiers and apply the receiver operating characteristic (ROC) curve fit method (described in [31]) to their outputs. This involves fitting the curve $\big{(}\varepsilon_{M_{1}}(t),\varepsilon_{M_{2}}(t)\big{)}$, where
@@
as determined by summing parton labels including weights, and $(f_{\text{Quarks}},f_{\text{Gluons}})=(1,0)$ by definition. Since we have observed that the quark and gluon distributions are sample independent to a good approximation in MC, we can attribute disagreement between the extractions and the MC predictions to some level mutual reducibility. In particular, the non-zero $\kappa$’s in Fig. 5a indicate that quarks and gluons in QCD MC are not irreducible in the output of the classifiers. The fully-supervised networks have reducibilities closest to zero, which is consistent with being the strongest quark/gluon discriminator in MC (classification metrics for each model will be presented in Sec. IV.3). In principle, one could use these estimates of $\kappa_{QG}$ and $\kappa_{GQ}$ to estimate quark fractions with these networks via Eq. 6.
Figure 5: Extracted reducibility factors for classifier outputs in each of the datasets. The red cross indicates the expectation for a mutually irreducible observable given fractions determined by MC parton labels.
@@
Classifier $\bm{\kappa_{QG}}$ $\bm{\kappa_{GQ}}$
Data CWoLa $0.169\pm 0.012$
MC CWoLa $0.157\pm 0.018$
Fully Supervised $0.249\pm 0.008$ $0.0711\pm 0.011$
Table 3: Quark/gluon reducibility factors of each classifier required to explain the measured $\kappa(\text{\footnotesize DY}|\text{\footnotesize QCD})$ and $\kappa(\text{\footnotesize QCD}|\text{\footnotesize DY})$ in Fig. 5b given the known MC quark fractions.
Figure 6: Quark fractions for DY and QCD simulated datasets (left column) and DoubleMu and Jet primary datasets (right column) according to jet topics extractions using each type of classifier compared to the parton level predictions. The top row assumes $\kappa_{QG}=\kappa_{GQ}=0$ (mutual irreducibility) for the network outputs whereas the bottom row assumes non-zero $\kappa$’s such that the correct fractions are recovered in MC.
@@
Figure 7: Distribution of Energy Flow Polynomials for quarks (top row) and gluons (bottom row) according to QCD MC and jet topics constructed from histograms of the primary datasets and TopicFlow. All topic distributions assume fractions determined from MC parton labels (Eq. 13). Bands show the mean $\pm$ one standard deviation over 10 flow models.
We now estimate the quark fractions of the mixed datasets. As discussed in Sec. II.2, if we assume the quark and gluon categories are mutually irreducible, then the fractions depend only on the reducibilities between $M_{1}$ and $M_{2}$ as in Eq. 4. Otherwise, the quark/gluon reducibility factors must be included per Eq. 6. The top row of Fig. 6 shows the fractions evaluated in the former fashion, assuming $\kappa_{QG}=\kappa_{GQ}=0$. In both the MC (6a) and primary (6b) datasets, the extractions are significantly different from the MC prediction shown by the red cross. When considering the primary datasets, the MC prediction is only a reference and the disagreement could be explained by mismodelling of quarks and gluons in simulation. Taking the Data CWoLa extraction as the representative estimate, we obtain quark fractions in the primary datasets of
@@
Table 4: Summary of our three estimates for the quark fractions of the DoubleMu and Jet datasets. $S$ (simulation) is taken directly from the sum of MC weights (Eq. 13). $T$ (topic) and $R$ (reducible) are extracted using Data CWoLa classifiers as described in Sec. IV.1. $T$ assumes quarks and gluons are mutually irreducible in the classifier score whereas $R$ assumes the same $\kappa_{QG}$ as MC (given in Table 3).
@@
While the above result assumes that the reducibility of $p_{Q}$ and $p_{G}$ in data is the same as in simulation, the result in Eq. 15 equally relies on assuming a value for $\kappa_{QG}$ (0 in that case). It is interesting that the ‘$\kappa$-corrected’ predictions in Fig. 6d do not reproduce the expected $f_{\mathrm{DoubleMu}},f_{\mathrm{Jet}}$ based on MC parton labels. This leads one to conclude that either $(f_{\mathrm{DoubleMu}},f_{\mathrm{Jet}})$ or $\kappa_{QG}$ is not well modelled in MC, assuming sample dependence does not playing a role. In the absence of another means to estimate the quark/gluon reducibilities in data, we will present results using each of the possible fractions. For clarity, we assign labels of $S$ (simulation), $T$ (topic) and $R$ (reducible) respectively to Eqs. 13, 15, and 16 as summarised in Table 4. Note that the question of reducibility is only relevant for extracting quark fractions, and the following analysis only requires that sample independence holds.
IV.2 Topic distributions
Here we present the results of TopicFlow trained on the DoubleMu and Jet datasets. Fig. 7 shows the results for quark and gluon topics, comparing both to ‘pure’ predictions from QCD simulation and the jet topic distributions formed by subtracting histograms of the mixed datasets according to Eq. 5. For best comparison with MC, all topic distributions (and TopicFlow loss) use fractions $S$. We also trained models assuming fractions $T$ and $R$, although the plots are qualitatively the same in each case. From each TopicFlow model, we generate 200k samples and the band shows the mean $\pm$ one standard deviation across ten models. Note that the samples are generated in the preprocessed basis, so the EFPs are recovered by inverting these transformations.
The first observation is that while the topic distributions constructed in MC mostly agree with the pure QCD predictions, neither MC distribution closely matches the topics constructed in data. The data-MC discrepancies are in fact larger than were displayed in Fig. 3 for the mixed simulation. This is despite using fractions $S$, which assume the same mixture proportions as MC. The TopicFlow networks generally provide a more precise model for quark/gluon topics in data than the simulation. We verified that the same results are found when comparing to DY simulation.
IV.3 Classification

### v2

@@ 2.2 Jet topics
The jet-topics framework Metodiev:2018ftz ; Komiske:2018vkc ; Komiske:2022vxg provides a method for extracting these fractions. In addition to sample independence of $p_{Q}$ and $p_{G}$ this method requires that one can measure an observable in which the distributions are mutually irreducible. This further assumption is equivalent to requiring that the support of $p_{Q}$ is not a subset of the support of $p_{G}$ and vice-versa. Under these conditions, Eq. 2 can be rewritten in terms of reducibility factors, defined as
$\kappa_{ij}\equiv\kappa(M_{i}|M_{j})=\underset{x}{\operatorname{min}}\,\frac{p_{M_{i}}(x)}{p_{M_{j}}(x)}.$ (3)
@@
Since the $\kappa$ factors are obtainable from samples of $M_{1}$ and $M_{2}$, their measurement provides a direct estimate for the quark fractions $f_{1},f_{2}$ and pure distributions that can be performed on data. Importantly, the extracted fractions apply to any observable, not only the $x$ with which the reducibility factors were determined. A number of methods for performing the minimisation of Eq. 3 are outlined in Komiske:2022vxg .
@@
Of course, categories defined in this manner are no longer completely data-driven. An alternative method for extracting mixture fractions in reducible observables is described in Zhu2023MixturePE .
2.3 TopicFlow
In previous work, some of us developed TopicFlow Dolan:2022ikg , with which one can train a generative model for the topic distributions $p_{Q}$ and $p_{G}$ given estimates of the quark fractions of two mixtures. The loss function of a deep generative model usually takes the form of an expectation value for some function $L$ (e.g. the negative log-likelihood in the case of normalising flows) over samples from the target distribution. While one cannot directly sample pure quark or gluon jets from experimental data, we can use Eq. 5 to rewrite the loss in terms of expectations over mixtures $M_{i}$. Taking the quark distribution as an example, one has
@@ 2.3 TopicFlow
In practice, training models with a loss that balances opposite-sign terms can harm the stability of the optimisation. In the case of the loss in Eq. 7, the network may discover that it can ‘cheat’ by ignoring the first term and trying to achieve poor loss on $M_{2}$, which is often trivial. In this work, we use an alternative loss that jointly trains quark and gluon topic models based on Eq. 1. Specifically, given a loss function $L$ that depends on the likelihood of the data under the model, we minimise [^3] For maximum likelihood training, the function $L(x)$ is simply $\log{x}$. One can of course consider other objective functions including those that depend on the likelihood only implicitly.
@@
While it is straightforward to apply Eq. 5 to histograms of two mixed datasets, generative models offer a number of benefits compared to binned approaches, particularly when $x$ is a high-dimensional observable. Firstly, generative models allow oversampling to smooth statistics over small datasets. In some cases, subject to the available data and the inductive bias of the model, deep generative models have been observed to contain more information than the data on which they were trained Butter:2020qhk ; Bieringer:2022cbs . A second advantage is the possibility of conditioning the generative model on auxiliary inputs such that distributions can be interpolated between measured points (e.g. Hallin:2021wme ; Algren:2023qnb ). The use case we will explore in this work, however, is generative classification where one constructs a quark/gluon likelihood ratio using individually modelled distributions.
In Dolan:2022ikg , TopicFlow was applied only to simulated data. However it is compatible with real data since the underlying model (a normalising flow) is unsupervised. This study therefore represents the first application of the method to real jets.
3 Training procedure and network architecture
@@
Fully supervised Quarks and gluons from QCD MC
@@
As a representation for the jets, we make use of Energy Flow Polynomials (EFPs) Komiske:2017aww , defined in Appendix B. EFPs form an over-complete basis for infrared and collinear safe (IRC) jet observables and are thus suited for use with real data, since physical detectors have finite energy/angle resolution. It was also observed in Bright-Thonney:2018mxq that IRC-safe observables exhibit smaller sample dependence. We consider connected EFPs with $\beta=\frac{1}{2}$ and number of edges $d\in[1,3\,]$.
Figure 4: Data/MC comparison for two Energy Flow Polynomials in LABEL:sub@efp_1_dijet, LABEL:sub@efp_2_dijet dijet events and LABEL:sub@efp_1_zjet, LABEL:sub@efp_2_zjet Z+Z+jet events.
Figure 4 shows histograms for two of the eight total EFPs in the basis, corresponding to the graphs
@@
The distributions exhibit small data-MC discrepancies, in particular systematically smaller EFP values in DY simulation compared to the DoubleMu primary dataset. As with the other substructure observables shown in Figure 3 above, the error bands are statistics-only.
@@
Figure 5: Histograms of the first two Energy Flow Polynomials (LABEL:sub@fig:efp_1_sample_dependence, LABEL:sub@fig:efp_2_sample_dependence) and classifier predictions (LABEL:sub@fig:mixed_preds, LABEL:sub@fig:mcmixed_preds, LABEL:sub@fig:pure_preds) for quarks and gluons in dijet and Z+Z+jet events from CMS simulation. The agreement among similar colours illustrates approximate sample independence in the datasets.
@@
For TopicFlow, we train continuous normalising flows with the Flow Matching objective of Lipman2022FlowMF . To adapt this loss according to Eq. 8, we replace the single regression network with a linear combination of ‘quark’ and ‘gluon’ networks as $v_{\theta}\to f_{i}v_{\theta_{Q}}+(1-f_{i})v_{\theta_{G}}$. [^5] Interestingly, for two vector fields $v_{Q}$ and $v_{G}$ respectively generating probability paths $p_{t,Q}$ and $p_{t,G}$ it does not follow that the combination $fv_{Q}+(1-f)v_{G}$ generates $fp_{t,Q}+(1-f)p_{t,G}$. Despite this, we find empirically that models trained in this way learn the expected topic distributions. For each regression network we use an MLP consisting of three 512-node layers with GELU activations and a skip connection between the first and third layer. Dropout is applied after each layer at a rate of 0.2. We use the Adam optimiser with an initial learning rate of $10^{-3}$ which is decayed by a factor 10 if there is no improvement for 20 epochs. Training is halted after 200 epochs, or once the learning rate has decayed below $10^{-6}$.
4 Results
4.1 Extracted quark fractions
@@
Here we investigate the quark/gluon content of each of the datasets using the jet topics. It is first necessary, however to assess whether or not the key assumptions of the framework are valid for our datasets. To validate the assumption of sample independence, we plot in Figure 5 the MC-matched quark and gluon distributions of two EFPs and the outputs of one of each type of classifier, both for DY and QCD simulation. In each observable, there are small but noticeable differences between the different MC datasets. However, the general shape of the distributions agree to a good approximation, with deviations mostly residing in the tails. In the EFP distributions, shown in the top row, the quark distributions are more similar across the Z+Z+jet and dijet samples compared to the gluon distributions. Interestingly, when considering the outputs of the two CWoLa models in Figs. 5c and 5d (which depend only on the 8 EFPs) the opposite trend is evident. Specifically, both of these models assign slightly higher scores to DY MC quarks than QCD MC quarks, as indicated by the uptick of purple squares in the DY/QCD ratio. The predictions of the CWoLa models exhibit the greatest sample dependence of the observables we consider, with deviations up to 25% in the worst regions. Meanwhile, the fully-supervised network has outputs that are relatively consistent across the MC samples.
@@
A fully-fledged measurement of any quantity requires a robust estimate of systematic uncertainties. In this essentially exploratory paper we have not attempted an analysis of these. All uncertainties presented are statistical. Consequently we refer to any quantities we derive in this paper as estimates and not as measurements.
@@
To extract reducibility factors and quark fractions in each dataset, we train classifiers and apply the receiver operating characteristic (ROC) curve fit method (described in Komiske:2022vxg ) to their outputs. This involves fitting the curve $\big{(}\varepsilon_{M_{1}}(t),\varepsilon_{M_{2}}(t)\big{)}$, where
@@
as determined by summing parton labels including weights, and $(f_{\text{Quarks}},f_{\text{Gluons}})=(1,0)$ by definition. Since we have observed that the quark and gluon distributions are sample independent to a good approximation in MC, we can attribute disagreement between the extractions and the MC predictions to some level of mutual reducibility. In particular, the non-zero $\kappa$’s in Fig. 6a indicate that quarks and gluons in QCD MC are not irreducible in the output of the classifiers. The fully-supervised networks have reducibilities closest to zero, which is consistent with being the strongest quark/gluon discriminator in MC (classification metrics for each model will be presented in Sec. 4.3). In principle, one could use these estimates of $\kappa_{QG}$ and $\kappa_{GQ}$ to estimate quark fractions with these networks via Eq. 6.
Figure 6: Extracted reducibility factors for classifier outputs in each of the datasets LABEL:sub@fig:pure_reducibilities Quark/Gluon simulation, LABEL:sub@fig:mcmixed_reducibilities DY/QCD simulation, LABEL:sub@fig:mixed_reducibilities DoubleMu/Jet primary. The red cross indicates the expectation for a mutually irreducible observable given fractions determined by MC parton labels.
@@
Classifier $\bm{\kappa_{QG}}$ $\bm{\kappa_{GQ}}$
Data CWoLa $0.169\pm 0.012$ -
MC CWoLa $0.157\pm 0.018$ -
Fully Supervised $249\pm 0.008 0.0$$0711\pm 0.011$
Table 3: Quark/gluon reducibility factors of each classifier required to explain the measured $\kappa(\text{\footnotesize DY}|\text{\footnotesize QCD})$ and $\kappa(\text{\footnotesize QCD}|\text{\footnotesize DY})$ in Fig. 6b given the known MC quark fractions. Uncertainties are calculated as the standard deviation over ten networks.
@@
Figure 7: Quark fractions for DY / QCD simulated datasets (LABEL:sub@fig:mcmixed_fractions, LABEL:sub@fig:mcmixed_fractions_corrected) and DoubleMu and Jet primary datasets (LABEL:sub@fig:mixed_fractions, LABEL:sub@fig:mixed_fractions_corrected) according to jet topics extractions using each type of classifier compared to the parton level predictions. The top row assumes $\kappa_{QG}=\kappa_{GQ}=0$ (mutual irreducibility) for the network outputs whereas the bottom row assumes non-zero $\kappa$’s such that the correct fractions are recovered in MC.
We now estimate the quark fractions of the mixed datasets. As discussed in Sec. 2.2, if we assume the quark and gluon categories are mutually irreducible, then the fractions depend only on the reducibilities between $M_{1}$ and $M_{2}$ as in Eq. 4. Otherwise, the quark/gluon reducibility factors must be included per Eq. 6. The top row of Fig. 7 shows the fractions evaluated in the former fashion, assuming $\kappa_{QG}=\kappa_{GQ}=0$. In both the MC (7a) and primary (7b) datasets, the extractions are different from the MC prediction shown by the red cross. When considering the primary datasets, the MC prediction is only a reference and the disagreement could be explained by mismodelling of quarks and gluons in simulation. Taking the Data CWoLa extraction as the representative estimate, we obtain quark fractions in the primary datasets of
@@
Table 4: Summary of our three estimates for the quark fractions of the DoubleMu and Jet datasets. $S$ (simulation) is taken directly from the sum of MC weights (Eq. 12). $T$ (topic) and $R$ (reducible) are extracted using Data CWoLa classifiers as described in Sec. 4.1. $T$ assumes quarks and gluons are mutually irreducible in the classifier score whereas $R$ assumes the same $\kappa_{QG}$ as MC (given in Table 3). The reported uncertainties are calculated as the standard deviation over the ten points.
@@
Again, uncertainties are the standard of ten points. While the above result assumes that the reducibility of $p_{Q}$ and $p_{G}$ in data is the same as in simulation, the result in Eq. 14 equally relies on assuming a value for $\kappa_{QG}$ (0 in that case). It is interesting that the $\kappa$-corrected’ predictions in Fig. 7d do not reproduce the expected $f_{\mathrm{DoubleMu}},f_{\mathrm{Jet}}$ based on MC parton labels. This leads one to conclude that either $(f_{\mathrm{DoubleMu}},f_{\mathrm{Jet}})$ or $\kappa_{QG}$ is not well modelled in MC, assuming sample dependence does not play a role. In the absence of another means to estimate the quark/gluon reducibilities in data, we will present results using each of the possible fractions. For clarity, we assign labels of $S$ (simulation), $T$ (topic) and $R$ (reducible) respectively to Eqs. 12, 14, and 15 as summarised in Table 4. Note that the question of reducibility is only relevant for extracting quark fractions, and the following analysis only requires that sample independence holds.
4.2 Topic distributions
Here we present the results of TopicFlow trained on the DoubleMu and Jet datasets.
Figure 8: Distribution of Energy Flow Polynomials for quarks (LABEL:sub@fig:d3_quark_efp_0_parton_topicflow, LABEL:sub@fig:d3_quark_efp_1_parton_topicflow) and gluons (LABEL:sub@fig:d3_gluon_efp_0_parton_topicflow, LABEL:sub@fig:d3_gluon_efp_1_parton_topicflow) according to QCD MC and jet topics constructed from histograms of the primary datasets and TopicFlow. All topic distributions assume fractions determined from MC parton labels (Eq. 12). Bands show the mean $\pm$ one standard deviation over 10 flow models.
Fig. 8 shows the results for quark and gluon topics, comparing both to ‘pure’ predictions from QCD simulation and the jet topic distributions formed by subtracting histograms of the mixed datasets according to Eq. 5. For best comparison with MC, all topic distributions (and TopicFlow loss) use fractions $S$. We also trained models assuming fractions $T$ and $R$, although the plots are qualitatively the same in each case. From each TopicFlow model, we generate 200k samples and the band shows the mean $\pm$ one standard deviation across ten models. Note that the samples are generated in the preprocessed basis, so the EFPs are recovered by inverting these transformations.
The first observation is that while the topic distributions constructed in MC mostly agree with the pure QCD predictions, neither MC distribution closely matches the topics constructed in data. The data-MC discrepancies are in fact larger than were displayed in Fig. 4 for the mixed simulation. This is despite using fractions $S$, which assume the same mixture proportions as MC. As discussed previously, it is possible that the data-MC difference is covered by systematics that we cannot quantify in this study. On the other hand, the TopicFlow networks closely follow the quark/gluon topics in data, showing no large biases in the bulk of the distributions. We verified that the same results are found when comparing to DY simulation.
4.3 Classification